# footnotetext: 2020 AMS Mathematics Subject Classification: 76B47, 35Q35footnotetext: Key words: Vortex stretching; Bi-rotational symmetry; vorticity; Biot–Savart law

### v1

@@ 1.2. Vorticity formulation of the bi-rotational solutions
The details of this computations are in the Section 2.
@@ 2. Computation of the vorticity and the stream function
In this section, we present computations of the vorticity and the stream function under the bi-rotational symmetry and no-swirl condition on the velocity (1.2) to show existences of the scalar vorticity $w$ and the scalar stream function $\psi$ satisfying (1.5).
@@ 2.1. Computation of the vorticity
(2.1) $u^{i}=\begin{cases}\frac{\partial r}{\partial x_{i}}u^{r}=\frac{x_{i}}{r}u^{r},&\quad i=1,\cdots,n+1,\\ \frac{\partial s}{\partial x_{i}}u^{s}=\frac{x_{i}}{s}u^{s},&\quad i=n+2,\cdots,n+m+2.\end{cases}$
@@
(2.2) $\begin{split}\omega^{i,j}&=\partial_{x_{j}}u^{i}-\partial_{x_{i}}u^{j}\\ &=\bigg[\frac{\partial r}{\partial x_{j}}\partial_{r}+\sum_{l=1}^{n}\frac{\partial\theta_{l}}{\partial x_{j}}\partial_{\theta_{l}}\bigg]\bigg(\frac{\partial r}{\partial x_{i}}u^{r}\bigg)-\bigg[\frac{\partial r}{\partial x_{i}}\partial_{r}+\sum_{l=1}^{n}\frac{\partial\theta_{l}}{\partial x_{i}}\partial_{\theta_{l}}\bigg]\bigg(\frac{\partial r}{\partial x_{j}}u^{r}\bigg)=0.\end{split}$
@@ 2.2. Computation of the stream function
(2.10) $\begin{split}\psi(r,s)&=\Delta^{-1}\omega^{1,n+2}({\bf x})\big|_{\theta_{1}=\cdots=\theta_{n}=\phi_{1}=\cdots=\phi_{m}=0}\\ &=c_{d}\int_{\mathbb{R}^{d}}\frac{1}{(r^{2}+\overline{r}^{2}-2ry_{1}+s^{2}+\overline{s}^{2}-2sy_{n+2})^{d/2-1}}\frac{y_{i}y_{j}}{\overline{r}\overline{s}}w(\overline{r},\overline{s})dy\\ &=c_{d}\iint_{\Pi}\iint_{[0,\pi]^{n+m}}\frac{\overline{r}^{n}\overline{s}^{m}\cos\overline{\theta}_{1}\cdots\cos\overline{\theta}_{n}\cos\overline{\phi}_{1}\cdots\cos\overline{\phi}_{m}}{(r^{2}+\overline{r}^{2}-2r\overline{r}\cos\overline{\theta}_{1}\cdots\cos\overline{\theta}_{n}+s^{2}+\overline{s}^{2}-2s\overline{s}\cos\overline{\phi}_{1}\cdots\cos\overline{\phi}_{m})^{d/2-1}}\cdot\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad w(\overline{r},\overline{s})d\overline{\theta}_{1}\cdots d\overline{\theta}_{n}d\overline{\phi}_{1}\cdots d\overline{\phi}_{m}d\overline{r}d\overline{s}.\end{split}$
Indeed, one can use the Biot–Savart law in the Cartesian coordinates (cf. [1])
(2.11) $u^{i}=\sum_{j=1}^{d}\partial_{x_{j}}\Psi^{i,j}$
to compute $u^{n+1}$ and $u^{n+m+2}$, and use the change of coordinates
$\frac{u^{r}}{r}=\frac{u^{n+1}}{x_{n+1}},\quad\frac{u^{s}}{s}=\frac{u^{n+m+2}}{x_{n+m+2}}$
@@
(2.12) $\begin{split}u^{r}&=\frac{r}{x_{n+1}}u^{n+1}=-\frac{r}{x_{n+1}}\bigg[\sin\theta_{1}\bigg(\partial_{s}\psi+\frac{m}{s}\psi\bigg)\bigg]=-\bigg(\partial_{s}\psi+\frac{m}{s}\psi\bigg)\\ u^{s}&=\frac{s}{x_{n+m+2}}u^{n+m+2}=\frac{s}{x_{n+m+2}}\bigg[\sin\phi_{1}\bigg(\partial_{r}\psi+\frac{n}{r}\psi\bigg)\bigg]=\partial_{r}\psi+\frac{n}{r}\psi.\end{split}$
@@ 3. Local well-posedness of the patch-type solution and blowup criterion
(3.4) $\begin{split}\|w(t)\|_{L^{\infty}(\mathbb{R}^{d})}&\leq R(t)^{n}S(t)^{m}\bigg\|\frac{w(t)}{r^{n}s^{m}}\bigg\|_{L^{\infty}(\mathbb{R}^{d})}=R(t)^{n}S(t)^{m}\bigg\|\frac{w_{0}}{r^{n}s^{m}}\bigg\|_{L^{\infty}(\mathbb{R}^{d})}\lesssim L(t)^{n+m}\bigg\|\frac{w_{0}}{r^{n}s^{m}}\bigg\|_{L^{\infty}(\mathbb{R}^{d})}.\end{split}$
@@
(3.5) $u^{i}(t,x)=\sum_{j=1}^{d}\partial_{x_{j}}\Psi^{i,j}(t,x),\quad i=1,\cdots,d,\quad t\geq 0,\quad x\in\mathbb{R}^{d},$
@@
$\dot{L}(t)\leq\|u(t)\|_{L^{\infty}(\mathbb{R}^{d})}\leq L(t)\|w(t)\|_{L^{\infty}(\mathbb{R}^{d})},$
@@
$\frac{d}{dt}(\ln L(t))\leq\|w(t)\|_{L^{\infty}(\mathbb{R}^{d})},$
@@
(3.23) $L(t)\leq L(0)\exp\bigg(\int_{0}^{t}\|w(\tau)\|_{L^{\infty}(\mathbb{R}^{d})}d\tau\bigg).$
@@ 4. Support growth
Now we plug in $w=\partial_{r}u^{s}-\partial_{s}u^{r}$ and use integration by parts to get
@@
In the third equality, we used the divergence-free condition (1.8). Finally, we use integration by parts one more time to obtain
@@
(4.9) $\begin{split}\partial_{s}\psi(r,0)&=c_{d}\iint_{\Pi}\iint_{[0,\pi/2]^{n+m}}\overline{r}^{n}\overline{s}^{m+1}\cos\overline{\theta}_{1}\cdots\cos\overline{\theta}_{n}(\cos\overline{\phi}_{1}\cdots\cos\overline{\phi}_{m})^{2}\cdot\\ &\qquad\bigg[\frac{1}{(r^{2}-2r\overline{r}\cos\overline{\theta}_{1}\cdots\cos\overline{\theta}_{n}+\overline{r}^{2}+\overline{s}^{2})^{d/2}}-\frac{1}{(r^{2}+2r\overline{r}\cos\overline{\theta}_{1}\cdots\cos\overline{\theta}_{n}+\overline{r}^{2}+\overline{s}^{2})^{d/2}}\bigg]\cdot\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad w(\overline{r},\overline{s})d\overline{\theta}_{1}\cdots d\overline{\theta}_{n}d\overline{\phi}_{1}\cdots d\overline{\phi}_{m}d\overline{r}d\overline{s}.\end{split}$
For simplicity, we denote $A^{\prime}:=r^{2}+\overline{r}^{2}+\overline{s}^{2}$ and $B^{\prime}:=2r\overline{r}\cos\overline{\theta}_{1}\cdots\cos\overline{\theta}_{n}$. Then using the inequality
@@
(4.11) $\begin{split}\partial_{s}\psi(r,0)&\gtrsim_{d}r\iint_{\Pi}\iint_{[0,\pi/2]^{n+m}}\frac{\overline{r}^{n+1}\overline{s}^{m+1}[\cos\overline{\theta}_{1}\cdots\cos\overline{\theta}_{n}\cos\overline{\phi}_{1}\cdots\cos\overline{\phi}_{m}]^{2}}{[(r+\overline{r})^{2}+\overline{s}^{2}]^{d/2+1}}\cdot\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad w(\overline{r},\overline{s})d\overline{\theta}_{1}\cdots d\overline{\theta}_{n}d\overline{\phi}_{1}\cdots d\overline{\phi}_{m}d\overline{r}d\overline{s}\\ &\gtrsim_{d}r\iint_{\Pi}\frac{\overline{r}^{n+1}\overline{s}^{m+1}}{[(r+\overline{r})^{2}+\overline{s}^{2}]^{d/2+1}}w(\overline{r},\overline{s})d\overline{r}d\overline{s}.\end{split}$

### v2

@@ 1.2. Vorticity formulation of the bi-rotational solutions
The details of these computations are in the Section 2.
@@ 2. Computation of the vorticity and the stream function
In this section, we present computations of the vorticity and the stream function under the bi-rotational symmetry and no-swirl condition on the velocity (1.2) to show existence of the scalar vorticity $w$ and the scalar stream function $\psi$ satisfying (1.5).
@@ 2.1. Computation of the vorticity
(2.1) $v^{i}=\begin{cases}\frac{\partial r}{\partial x_{i}}u^{r}=\frac{x_{i}}{r}u^{r},&\quad i=1,\cdots,n+1,\\ \frac{\partial s}{\partial x_{i}}u^{s}=\frac{x_{i}}{s}u^{s},&\quad i=n+2,\cdots,n+m+2.\end{cases}$
@@
(2.2) $\begin{split}\omega^{i,j}&=\partial_{x_{j}}v^{i}-\partial_{x_{i}}v^{j}\\ &=\bigg[\frac{\partial r}{\partial x_{j}}\partial_{r}+\sum_{l=1}^{n}\frac{\partial\theta_{l}}{\partial x_{j}}\partial_{\theta_{l}}\bigg]\bigg(\frac{\partial r}{\partial x_{i}}u^{r}\bigg)-\bigg[\frac{\partial r}{\partial x_{i}}\partial_{r}+\sum_{l=1}^{n}\frac{\partial\theta_{l}}{\partial x_{i}}\partial_{\theta_{l}}\bigg]\bigg(\frac{\partial r}{\partial x_{j}}u^{r}\bigg)=0.\end{split}$
@@ 2.2. Computation of the stream function
(2.10) $\begin{split}&\psi(r,s)=-\Delta^{-1}\omega^{1,n+2}({\bf x})\big|_{\theta_{1}=\cdots=\theta_{n}=\phi_{1}=\cdots=\phi_{m}=0}\\ &\;=c_{d}\int_{\mathbb{R}^{d}}\frac{1}{(r^{2}+\overline{r}^{2}-2ry_{1}+s^{2}+\overline{s}^{2}-2sy_{n+2})^{d/2-1}}\frac{y_{1}y_{n+2}}{\overline{r}\overline{s}}w(\overline{r},\overline{s})dy\\ &\;=c_{d}\iint_{\Pi}\int_{0}^{2\pi}\int_{[0,\pi]^{m-1}}\int_{0}^{2\pi}\int_{[0,\pi]^{n-1}}\frac{1}{(r^{2}+\overline{r}^{2}-2r\overline{r}\cos\overline{\theta}_{1}+s^{2}+\overline{s}^{2}-2s\overline{s}\cos\overline{\phi}_{1})^{d/2-1}}\\ &\quad\cdot\cos\overline{\theta}_{1}\cos\overline{\phi}_{1}w(\overline{r},\overline{s})\bigg(\overline{r}^{n}\overline{s}^{m}\prod_{i=1}^{n-1}\sin^{n-i}\overline{\theta}_{i}\prod_{j=1}^{m-1}\sin^{m-j}\overline{\phi}_{j}\bigg)d\overline{\theta}_{1}\cdots d\overline{\theta}_{n-1}d\overline{\theta}_{n}d\overline{\phi}_{1}\cdots d\overline{\phi}_{m-1}d\overline{\phi}_{m}d\overline{r}d\overline{s}\\ &\;=c_{d}\iint_{\Pi}\int_{0}^{\pi}\int_{0}^{\pi}\frac{\overline{r}^{n}\overline{s}^{m}\sin^{n-1}\overline{\theta}_{1}\cos\overline{\theta}_{1}\sin^{m-1}\overline{\phi}_{1}\cos\overline{\phi}_{1}}{(r^{2}+\overline{r}^{2}-2r\overline{r}\cos\overline{\theta}_{1}+s^{2}+\overline{s}^{2}-2s\overline{s}\cos\overline{\phi}_{1})^{d/2-1}}w(\overline{r},\overline{s})d\overline{\theta}_{1}d\overline{\phi}_{1}d\overline{r}d\overline{s},\end{split}$
where $c_{d}>0$ denotes a constant that depends only on $d$ which may change line by line. Indeed, one can use the Biot–Savart law in the Cartesian coordinates (cf. [1])
(2.11) $v^{i}=\sum_{j=1}^{d}\partial_{x_{j}}\Psi^{i,j}$
to compute $v^{1}$ and $v^{n+2}$, and use the change of coordinates
(2.12) $\frac{u^{r}}{r}=\frac{v^{1}}{x_{1}},\quad\frac{u^{s}}{s}=\frac{v^{n+2}}{x_{n+2}},$
@@
(2.13) $\begin{split}u^{r}&=\frac{r}{x_{1}}v^{1}=-\frac{r}{x_{1}}\bigg[\cos\theta_{1}\bigg(\partial_{s}\psi+\frac{m}{s}\psi\bigg)\bigg]=-\bigg(\partial_{s}\psi+\frac{m}{s}\psi\bigg),\\ u^{s}&=\frac{s}{x_{n+2}}v^{n+2}=\frac{s}{x_{n+2}}\bigg[\cos\phi_{1}\bigg(\partial_{r}\psi+\frac{n}{r}\psi\bigg)\bigg]=\partial_{r}\psi+\frac{n}{r}\psi.\end{split}$
@@ 3. Local well-posedness of the patch-type solution and blowup criterion
(3.4) $\begin{split}\|w(t)\|_{L^{\infty}(\mathbb{R}^{d})}&\leq R(t)^{n}S(t)^{m}\bigg\|\frac{w(t)}{r^{n}s^{m}}\bigg\|_{L^{\infty}(\mathbb{R}^{d})}=R(t)^{n}S(t)^{m}\bigg\|\frac{w_{0}}{r^{n}s^{m}}\bigg\|_{L^{\infty}(\mathbb{R}^{d})}\leq L(t)^{n+m}\bigg\|\frac{w_{0}}{r^{n}s^{m}}\bigg\|_{L^{\infty}(\mathbb{R}^{d})}.\end{split}$
@@
(3.5) $v^{i}(t,x)=\sum_{j=1}^{d}\partial_{x_{j}}\Psi^{i,j}(t,x),\quad i=1,\cdots,d,\quad t\geq 0,\quad x\in\mathbb{R}^{d},$
@@
$\dot{L}(t)\leq\|u(t)\|_{L^{\infty}(\mathbb{R}^{d})}\lesssim_{d}L(t)\|w(t)\|_{L^{\infty}(\mathbb{R}^{d})},$
@@
$\frac{d}{dt}(\ln L(t))\lesssim_{d}\|w(t)\|_{L^{\infty}(\mathbb{R}^{d})},$
@@
(3.23) $L(t)\lesssim_{d}L(0)\exp\bigg(\int_{0}^{t}\|w(\tau)\|_{L^{\infty}(\mathbb{R}^{d})}d\tau\bigg).$
@@ 4. Support growth
The divergence-free condition (1.8) was used in the second equality. Now we plug in $w=\partial_{r}u^{s}-\partial_{s}u^{r}$ and use integration by parts to get
@@
In the third equality, we used the divergence-free condition (1.8) once more. Finally, we use integration by parts one more time to obtain
@@
(4.9) $\begin{split}&\partial_{s}\psi(r,0)=c_{d}\iint_{\Pi}\int_{0}^{\pi}\int_{0}^{\pi}\frac{\overline{r}^{n}\overline{s}^{m+1}\sin^{n-1}\overline{\theta}_{1}\cos\overline{\theta}_{1}\sin^{m-1}\overline{\phi}_{1}\cos^{2}\overline{\phi}_{1}}{(r^{2}+\overline{r}^{2}-2r\overline{r}\cos\overline{\theta}_{1}+\overline{s}^{2})^{d/2-1}}w(\overline{r},\overline{s})d\overline{\theta}_{1}d\overline{\phi}_{1}d\overline{r}d\overline{s}\\ &\quad=c_{d}\iint_{\Pi}\int_{0}^{\pi/2}\int_{0}^{\pi/2}\overline{r}^{n}\overline{s}^{m+1}\sin^{n-1}\overline{\theta}_{1}\cos\overline{\theta}_{1}\sin^{m-1}\overline{\phi}_{1}\cos^{2}\overline{\phi}_{1}\\ &\qquad\cdot\bigg[\frac{1}{(r^{2}+\overline{r}^{2}-2r\overline{r}\cos\overline{\theta}_{1}+\overline{s}^{2})^{d/2-1}}-\frac{1}{(r^{2}+\overline{r}^{2}+2r\overline{r}\cos\overline{\theta}_{1}+\overline{s}^{2})^{d/2-1}}\bigg]w(\overline{r},\overline{s})d\overline{\theta}_{1}d\overline{\phi}_{1}d\overline{r}d\overline{s}.\end{split}$
For simplicity, we denote $A^{\prime}:=r^{2}+\overline{r}^{2}+\overline{s}^{2}$ and $B^{\prime}=2r\overline{r}\cos\overline{\theta}_{1}$. Then using the inequality
@@
(4.11) $\begin{split}\partial_{s}\psi(r,0)&\gtrsim_{d}r\iint_{\Pi}\int_{0}^{\pi/2}\int_{0}^{\pi/2}\frac{\overline{r}^{n+1}\overline{s}^{m+1}\sin^{n-1}\overline{\theta}_{1}\cos^{2}\overline{\theta}_{1}\sin^{m-1}\overline{\phi}_{1}\cos^{2}\overline{\phi}_{1}}{[(r+\overline{r})^{2}+\overline{s}^{2}]^{d/2+1}}w(\overline{r},\overline{s})d\overline{\theta}_{1}d\overline{\phi}_{1}d\overline{r}d\overline{s}\\ &\gtrsim_{d}r\iint_{\Pi}\frac{\overline{r}^{n+1}\overline{s}^{m+1}}{[(r+\overline{r})^{2}+\overline{s}^{2}]^{d/2+1}}w(\overline{r},\overline{s})d\overline{r}d\overline{s}.\end{split}$